OpenStudy (anonymous):

The Value of an uncirculated Mint State 65' 1950 Jefferson nickel minted in Denver is 3/2 the value of a 1945 nickel minted in Philadelphia in simular condition. Together the total value of the two coins is 35$. What is the value of each coin?

OpenStudy (turingtest):

want the process or answer?

OpenStudy (anonymous):

Both please

OpenStudy (anonymous):

If it's not too much trouble, I can't seem to figure out how to do this problem

OpenStudy (turingtest):

x= value of the Denver coin y= value of the Philadelphia coin y=(3/2)x x+y=35 (total value of the two coins) sub the expression for y into the previous equation: x+(3/2)x=35 2x+3x=70 5x=70 x=14 (the Denver coin is worth $14) y=(3/2)x=(3/2)14=21 (the Philly coin is worth $21) get it?

OpenStudy (anonymous):

yes I do, thank you very much, I really appreciate it

OpenStudy (turingtest):

how 'bout a medal? ;)